OpenStudy (anonymous):

sqrt 5x-1 = sqrt 4x+9

OpenStudy (anonymous):

x=10

OpenStudy (anonymous):

( sqrt(5x-1))^2 = (sqrt(4x+9)^2 5x-1 = 4x+9 x=10

OpenStudy (anonymous):

I didn't expect an answer so fast.. thank you, guys! You guys' are awesome! xoxo